OpenStudy (lexy148):

the listing agent for a house receives 20% of the agency fee as commission when a sale occurs even if another realtor sells the house. If a house sells for $250,000 and the agency receives 6% of the selling price, how much will the listing agent be paid?

OpenStudy (anonymous):

the agency gets 6% of selling price so, agency gets commission = 6% of 250,000 = 6/100* 250,000 = 15000 as per question, listing agent gets 20% of agency commission so gets = 20% of 15000 = 20/100 *15000 = $3000

OpenStudy (lexy148):

thank you

OpenStudy (anonymous):

my pleasure